space 1 2 1 3 1 2 4 3
P(space) 0.125 0.125 0.125 0.125 0.125 0.125 0.125 0.125

What is the theoretical probability of moving 4 or less spaces in 3 spins if each spin is independent?
Answer #1

The theoretical probability of moving 4 or less spaces in 1 spin = p = P(space = 1) + P(space = 2) + P(space = 3) + P(space = 4) = 0.125 + 0.125 + 0.125 + 0.125 = 0.5

the theoretical probability of moving 4 or less spaces in 3 spins if each spin is independent = p3 = 0.53 = 0.125

So,

Answer is:

0.125